# Non-stationary Spatio-Temporal Modeling Using the Stochastic Advection-Diffusion Equation

### v1

@@ A.3 Spatial discretization
$\phantom{=}-\Delta t\oint_{\partial E_{i,j}}\left[\nabla u^{n+1}(\bm{s})\cdot(\mathbf{H}(\bm{s}){\bm{n}}(\bm{s}))+u^{n+1}(\bm{s})(\bm{\omega}(\bm{s})\cdot{\bm{n}}(\bm{s})\right]\mathrm{d}A$
@@
The second line is subtracting the net flow leaving the grid cell. For approximating the flow arising from diffusion we refer to Fuglstad et al., 2015a , and we detail the approximation of the flow arising from advection in the next subsection. Note that the zero flow boundary condition means that the line integral on the second line is 0 on faces that belong to the boundary, which is desired as there should be zero flow out of the boundary.
@@ A.4 Approximating flow arising from advection
$\int_{A_{i,j}^{\text{Dir}}}u(\bm{s})\left(\bm{\omega}(\bm{s})\cdot{\bm{n}}(\bm{s})\right)\mathrm{d}A\approx$ $u(\bm{s}_{i,j})^{\text{Dir}}\left(\bm{\omega}(\bm{s}_{i,j}^{\text{Dir}})\cdot{\bm{n}}(\bm{s}_{i,j}^{\text{Dir}})\right)\int_{A_{i,j}^{\text{Dir}}}\mathrm{d}A$
@@
We use an upwind scheme, where we take the direction of the flow into account. When considering the flow over a given face, we only consider the upwind cell according to the advection vector field at the centre of the face. E.g., if we are considering the right face of a grid cell and the flow goes to the left, we use the value of the field in the grid cell to the right. This is known to help with numerical stability and is a common scheme in the finite volume method.
@@
We vectorize the above expressions and define the vector of all spatial locations for timestep $n+$ as
@@
for $n=0,\ldots,T-1$. We introduce the advection matrix $\mathbf{A}_{\bm{\omega}}$ such that $\mathbf{A}_{\bm{\omega}}{\bm{u}}^{n+1}$ gives the advection term in Equation (S3) using the above approximations. Thus $\mathbf{A}_{\bm{\omega}}$ is a $MN\times MN$ matrix, where the local solution of the advection term for location $(i,j)$ is collected in the $(jM+i)$th row of the matrix. In the following equations, we show the non-zero elemests. For the diagonal elements we have
@@ A.5 The fully descritized SPDE and the precision matrix
where ${\bm{u}}^{0}\sim\mathcal{N}(\bm{0},\mathbf{Q}_{0})$ arising from approximating $\text{WM}(\kappa_{\mathrm{I}}(\cdot),\mathbf{H}_{\mathrm{I}}(\cdot))$. However, note that computationally one would not compute the dense matrices $\mathbf{G}$ and $\mathbf{E}$ when using this equation for simulation.
@@
$\pi\left(\begin{bmatrix}{\bm{u}}^{0}\\ \bm{\epsilon}\end{bmatrix}\right)\propto\exp\left(-\frac{1}{2}\begin{bmatrix}({\bm{u}}^{0})^{\mathrm{T}}&\bm{\epsilon}^{\mathrm{T}}\end{bmatrix}\begin{bmatrix}\mathbf{Q}_{0}&0&\dots&0\\ 0&\mathbf{I}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&\mathbf{I}\\ \end{bmatrix}\begin{bmatrix}{\bm{u}}^{0}\\ \bm{\epsilon}\end{bmatrix}\right).$
@@
$=$ $\exp\left(-\frac{1}{2}{\bm{u}}^{\top}\mathbf{R}^{-{\top}}\begin{bmatrix}\mathbf{Q}_{0}&0&\dots&0\\ 0&\mathbf{I}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&\mathbf{I}\\ \end{bmatrix}\mathbf{R}^{-1}{\bm{u}}\right).$
@@
$\mathbf{Q}=\begin{bmatrix}\mathbf{Q}_{0}+\mathbf{G}^{\top}\mathbf{F}^{-1}\mathbf{G}&-\mathbf{G}^{\top}\mathbf{F}^{-1}&0&\dots&0\\ -\mathbf{F}^{-1}\mathbf{G}&\mathbf{F}^{-1}+\mathbf{G}^{\top}\mathbf{F}^{-1}\mathbf{G}&-\mathbf{G}^{\top}\mathbf{F}^{-1}&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&-\mathbf{F}^{-1}\mathbf{G}&\mathbf{F}^{-1}+\mathbf{G}^{\top}\mathbf{F}^{-1}\mathbf{G}&-\mathbf{G}^{\top}\mathbf{F}^{-1}\\ 0&\dots&0&-\mathbf{F}^{-1}\mathbf{G}&\mathbf{F}^{-1}\end{bmatrix}.$
Appendix B. Code implementations
The models described in this work are implemented in the Python package currently named spdepy. The package is available at https://github.com/berild/spdepy and is under development. Currently, the package is not available on PyPI, but can be installed using poetry.
@@ Appendix B. Code implementations
Available models in the package are the advection-diffusion SPDE, the separable space-time model, and the Whittle-Matérn SPDE. The user is free to specify if the parameters are spatially varying or not, and many different combinations are allowed. For example, non-stationary anisotropic spatial fields or simply stationary isotropic fields, and spatially varying advection and diffusion with a simple stationary isotropic initial field or simple constant advection and spatially varying diffusion with a complex non-stationary anisotropic initial field. It is also possible to set the advection field as a covariate.

### v2

@@ A.3 Spatial discretization
$\phantom{=}-\Delta t\oint_{\partial E_{i,j}}\left[-\nabla u^{n+1}(\bm{s})\cdot\{\mathbf{H}(\bm{s}){\bm{n}}(\bm{s})\}+u^{n+1}(\bm{s})\{\bm{\omega}(\bm{s})\cdot{\bm{n}}(\bm{s})\}\right]\mathrm{d}A$
@@
The second line is subtracting the net flow leaving the grid cell. For approximating the flow arising from diffusion we refer to Fuglstad et al., 2015a , and we detail the approximation of the flow arising from advection in the next subsection.
@@ A.4 Approximating flow arising from advection
$\int_{A_{i,j}^{\text{Dir}}}u(\bm{s})\left(\bm{\omega}(\bm{s})\cdot{\bm{n}}(\bm{s})\right)\mathrm{d}A\approx$ $u(\bm{s}_{i,j}^{\text{Dir}})\left(\bm{\omega}(\bm{s}_{i,j}^{\text{Dir}})\cdot{\bm{n}}(\bm{s}_{i,j}^{\text{Dir}})\right)\int_{A_{i,j}^{\text{Dir}}}\mathrm{d}A$
@@
We use an upwind scheme (Eymard et al.,, 2000, Section 20.3), where we take the direction of the flow into account. When considering the flow over a given face, we only consider the upwind cell according to the advection vector field at the centre of the face. E.g., if we are considering the right face of a grid cell and the flow goes to the left, we use the value of the field in the grid cell to the right. This is known to help with numerical stability and is a common scheme in the finite volume method.
@@
We vectorize the above expressions and define the vector of all spatial locations for timestep $n$ as
@@
for $n=0,\ldots,T-1$. We introduce the advection matrix $\mathbf{A}_{\bm{\omega}}$ such that $\mathbf{A}_{\bm{\omega}}{\bm{u}}^{n+1}$ gives the advection term in Equation (S3) using the above approximations. Thus $\mathbf{A}_{\bm{\omega}}$ is a $MN\times MN$ matrix, where the local solution of the advection term for location $(i,j)$ is collected in the $(jM+i)$th row of the matrix. In the following equations, we show the non-zero elements. For the diagonal elements we have
@@ A.5 The fully descritized SPDE and the precision matrix
where ${\bm{u}}^{0}\sim\mathcal{N}(\bm{0},\mathbf{Q}_{\mathrm{I}}^{-1})$ arising from approximating $\text{WM}(\kappa_{\mathrm{I}}(\cdot),\mathbf{H}_{\mathrm{I}}(\cdot))$. However, note that computationally one would not compute the dense matrices $\mathbf{G}$ and $\mathbf{E}$ when using this equation for simulation.
@@
$\pi\left(\begin{bmatrix}{\bm{u}}^{0}\\ \bm{\epsilon}\end{bmatrix}\right)\propto\exp\left(-\frac{1}{2}\begin{bmatrix}({\bm{u}}^{0})^{\mathrm{T}}&\bm{\epsilon}^{\mathrm{T}}\end{bmatrix}\begin{bmatrix}\mathbf{Q}_{\mathrm{I}}^{-1}&0&\dots&0\\ 0&\mathbf{I}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&\mathbf{I}\\ \end{bmatrix}\begin{bmatrix}{\bm{u}}^{0}\\ \bm{\epsilon}\end{bmatrix}\right).$
@@
$=$ $\exp\left(-\frac{1}{2}{\bm{u}}^{\top}\mathbf{R}^{-{\top}}\begin{bmatrix}\mathbf{Q}_{\mathrm{I}}&0&\dots&0\\ 0&\mathbf{I}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&\mathbf{I}\\ \end{bmatrix}\mathbf{R}^{-1}{\bm{u}}\right).$
@@
$\mathbf{Q}=\begin{bmatrix}\mathbf{Q}_{\mathrm{I}}+\mathbf{G}^{\top}\mathbf{F}^{-1}\mathbf{G}&-\mathbf{G}^{\top}\mathbf{F}^{-1}&0&\dots&0\\ -\mathbf{F}^{-1}\mathbf{G}&\mathbf{F}^{-1}+\mathbf{G}^{\top}\mathbf{F}^{-1}\mathbf{G}&-\mathbf{G}^{\top}\mathbf{F}^{-1}&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&-\mathbf{F}^{-1}\mathbf{G}&\mathbf{F}^{-1}+\mathbf{G}^{\top}\mathbf{F}^{-1}\mathbf{G}&-\mathbf{G}^{\top}\mathbf{F}^{-1}\\ 0&\dots&0&-\mathbf{F}^{-1}\mathbf{G}&\mathbf{F}^{-1}\end{bmatrix}.$
Appendix B Code implementations
The models described in this work are implemented in a Python package named spdepy. The package is available at https://github.com/berild/spdepy and is under development. Currently, the package is not available on PyPI, but can be installed using poetry.
@@ Appendix B Code implementations
Available SPDEs in the package include the advection-diffusion SPDE, the separable space-time model, and the Whittle-Matérn SPDE. Users have the flexibility to specify whether the parameters are spatially varying or constant, allowing for various combinations. For instance, spatial models can be stationary or non-stationary, as well as isotropic or anisotropic. Additionally, spatio-temporal models can incorporate constant or spatially varying advection and diffusion, or any combination of these factors. It is also possible to treat advection as a covariate. Users can input any initial condition, provided it aligns with one of the aforementioned models included in the package.